OpenStudy (anonymous):

What is the result when a number that is 20% less than 48 is reduced by 15%

OpenStudy (zzr0ck3r):

(48-0.20*48)-.15(48-0.20*48)

OpenStudy (zzr0ck3r):

I think

OpenStudy (mathstudent55):

A number that is 20% less than 48 is 80% of 48. A number that is reduced by 15% is 85% of what is used to be. Just multiply 0.85 * 0.8 * 48

OpenStudy (zzr0ck3r):

his way is much better^^^^

OpenStudy (zzr0ck3r):

his/her...

OpenStudy (anonymous):

That doesn't help me :(

OpenStudy (mathstudent55):

\( (48-0.20 \times 48)- 0.15(48-0.20 \times 48) \) \(= 48(1-0.20 \times 1)- 0.15 \times 48 (1-0.20 \times 1) \) \(= 48(0.8)- 0.15 \times 48 (0.8 \times 1) \) \(= 48(0.8)- 0.15 \times 48 (0.8) \) \(= 48(0.8)(1 - 0.15 \times 1) \) \(= 48(0.8)(1 - 0.15) \) \(= 48(0.8)(0.85) \) \(= 0.85 \times 0.8 \times 48 \) \(= 32.64\)

OpenStudy (anonymous):

Still don't get it, but whatever.

OpenStudy (mathstudent55):

@Quinn33 I'll explain.

OpenStudy (mathstudent55):

100% means a complete amount. It means 100 out of 100. If you get 100% on a test, that means you got every question correct. Ok so far?

OpenStudy (anonymous):

Yes

OpenStudy (mathstudent55):

If you take a number and you reduce it a certain percentage, such as reduce a number by 20%, that means instead of having 100% of the number, you want only 100% - 20% which is 80% of the number.

OpenStudy (mathstudent55):

In other words, the number that is 20% less than 48 is 80% of 48. To find a percentage of a number, just multiply the percentage (written as a decimal) by the number. 80% of 48 = 80% * 48 = 0.8 * 48

OpenStudy (mathstudent55):

Now we have the number that is 20% less than 48, it's 0.8 * 48. Now we need to reduce that number by 15%.

OpenStudy (anonymous):

the number that is 20% less than 48 is: \(\Large a=48-48\times \frac{20}{100}=38.4\) And \(\Large a\) is reduced by 15%, \(\Large \Rightarrow a-\frac{15}{100}a=38.4-38.4\times \frac{15}{100}=32.64\)

OpenStudy (mathstudent55):

Reducing 15% means 100% = 15% = 85%. When you reduce a number 15%, you are taking 85% of the number. How do you reduce 0.8 * 48 by 15%? You multiply 0.8 * 48 by 85%. 85% + 0.8 * 48 = 0.85 * 0.8 * 48 Just multiply 0.85 * 0.8 * 478 to find the answer.